# Predicting Case Suffixes With Activity Start and End Times: A Sweep-Line Based Approach

### v1

@@ 27 $\text{WIP}_{c,k}\leftarrow$ TotalActiveInstances;
28  $r_{c,k}\leftarrow r_{c,k}\oplus(\text{WIP}_{c,k})$;
@@
30  $r_{c,k}\leftarrow r_{c,k}\oplus(\text{Utilization}_{c,k})$;
@@
32  Initialize RecentCaseStarts $\leftarrow\emptyset$;
@@ 34     Let $T^{\text{start}}_{c^{\prime}}$ be the first start timestamp in $\sigma_{c^{\prime}}$;
35     if $T^{\text{start}}_{c^{\prime}}\geq T^{\text{eval}}-\tau$ then
@@ 36        Add $c^{\prime}$ to RecentCaseStarts;
40  $r_{c,k}\leftarrow r_{c,k}\oplus(\lambda_{c,k})$;
41
Algorithm 1 Enhancing an activity instance trace $\sigma_{c}$ with Inter & Intra-case Features
3.1.2 Intra-Case Features:
@@
We extract features from each trace prefix of a trace $\sigma_{c}$. Given a trace $\sigma_{c}=\langle r_{c,1},\dots,r_{c,n}\rangle$ capturing the execution of case $c$, and given an index $1\leq k\leq n$, the trace prefix of case $c$ at index $k$, denoted $\text{$\mathcal{P}$}(c,k)$, is the sequence $\langle r_{c,1},\dots,r_{c,k}\rangle$. For example, in Table 1, the prefix of case 1 at index 2 $\text{$\mathcal{P}$}(1,2)=\langle(\text{Received Query},08{:}00,08{:}05),(\text{Assigned to Rep},08{:}10,08{:}15)\rangle$.
@@
For each activity instance $r_{c,k}$ = $(c,a_{c,k},T^{\text{start}}_{c,k},T^{\text{end}}_{c,k})$ in a prefix, we construct an enhanced activity instance $(c,a_{c,k},T^{\text{start}}_{c,k},T^{\text{end}}_{c,k},\delta^{\text{start}}_{c,k},\delta^{\text{proc}}_{c,k})$, where $\delta^{\text{start}}_{c,k}=T^{\text{start}}_{c,k}-T^{\text{start}}_{c,k-1}$ denotes the inter-start time, and $\delta^{\text{proc}}_{c,k}=T^{\text{end}}_{c,k}-T^{\text{start}}_{c,k}$ is the processing time. For the first activity, we define $\delta^{\text{start}}_{c,1}=0$. The procedure for extracting vector $v$ is implemented in lines (1-8) of Alg. 1. The algorithm iterates over each trace $\sigma_{c}\in\mathcal{L}$, where $\mathcal{L}$ is the activity instance log grouped by case. For each trace prefix $\text{$\mathcal{P}$}\subseteq\sigma_{c}$, it processes every activity instance, computes the associated intra-case features, and concatenates these features with the tuple representing activity instance $r_{c,i}$ with these features (line 9). [^1] Operator $\oplus$ denotes tuple concatenation.
3.1.3 Inter-Case Features:
Inter-case features capture dependencies across multiple cases, influencing resource contention and case concurrency. Unlike intra-case features, they help predict waiting times and delays. These features are crucial for modeling process dynamics, particularly in resource-constrained systems. Prior research [8] highlights the importance of inter-case features such as work in progress (WIP) and process load (e.g., resource utilization) in capturing dynamic process behavior. Alg. 1 explains the extraction of inter-case features for a given $\mathcal{P}$. It computes three inter-case features: WIP, resource utilization, and the arrival rate of new cases (lines 10-29).
According to Little’s Law [5], the average cycle time ($T_{c}$) of a process is determined by the Work in Progress (WIP) and arrival rate ($\lambda$) via $WIP=\lambda\times T_{c}$. In human-centric processes, cycle time is often dominated by waiting time. We, therefore, hypothesize that features correlated with cycle time can improve waiting time prediction. To compute WIP at time $T^{\text{start}}_{c,k}$, Alg. 1 counts active activity instances in $\mathcal{L}{}$ (lines 11-17). It also estimates the arrival rate as $\lambda=\frac{\text{Total cases}}{\text{Time window}}$ (lines 22-28), using a recent time window (default: 20% of log duration). The set of recent cases is updated dynamically (line 27). Higher WIP and $\lambda$ indicate system congestion, leading to longer expected waiting times.
Queuing theory suggests that waiting times are influenced by resource utilization, the ratio of resource demand to availability [2]. When demand exceeds capacity, utilization reaches 100%, causing delays. Lower demand reduces utilization. Alg. 1 computes utilization by counting active instances per activity at a given time (lines 16-20), helping predictive models account for resource load and capacity. For each trace prefix $\text{$\mathcal{P}$}\subseteq\sigma_{c}$, Alg. 1 computes inter-case features and augments $r_{c,k}$ accordingly.
3.1.4 Sequence Encoding and Construction:

### v2

@@ 36 $\text{WIP}_{c,k}\leftarrow$ TotalActiveInstances;
37 $e_{c,k}\leftarrow e_{c,k}\oplus(\text{WIP}_{c,k})$;
38
@@ 37 $e_{c,k}\leftarrow e_{c,k}\oplus(\text{WIP}_{c,k})$;
40 $e_{c,k}\leftarrow e_{c,k}\oplus(\text{Utilization}_{c,k})$;
41
@@ 40 $e_{c,k}\leftarrow e_{c,k}\oplus(\text{Utilization}_{c,k})$;
43 RecentCaseStarts $\leftarrow\emptyset$;
44
@@ 46      Let $T^{\text{start}}_{c^{\prime}}$ be the first start timestamp in $\sigma_{c^{\prime}}$;
47    if $T^{\text{start}}_{c^{\prime}}\geq T^{\text{current}}-\tau$ then
@@ 48         Add $c^{\prime}$ to RecentCaseStarts;
55 $e_{c,k}\leftarrow e_{c,k}\oplus(\lambda_{c,k})$;
56
 // --- Resource Availability Features ---
57
58 $\mathcal{R}^{a_{k}}\leftarrow\mathcal{R}(a_{c,k})$;
59
60 foreach $r\in\mathcal{R}^{a_{k}}$ do
61    $T^{r}_{\text{next\_on\_duty}}\leftarrow\textsc{NextOnDuty}(C_{r},T^{\text{current}})$;
62
63    $T^{r}_{\text{since\_last\_on\_duty}}\leftarrow\textsc{SinceLastOnDuty}(C_{r},T^{\text{current}})$;
64
65    $T^{r}_{\text{next\_off\_duty}}\leftarrow\textsc{NextOffDuty}(C_{r},T^{\text{current}})$;
66
67    $T^{r}_{\text{since\_last\_off\_duty}}\leftarrow\textsc{SinceLastOffDuty}(C_{r},T^{\text{current}})$;
68
69   end foreach
70
71 foreach $f\in\{T_{\text{next\_on\_duty}},T_{\text{since\_last\_on\_duty}},T_{\text{next\_off\_duty}},T_{\text{since\_last\_off\_duty}}\}$ do
72    $\bar{f}(T^{\text{current}})\leftarrow\min_{r\in\mathcal{R}^{a_{k}}}f(r,T^{\text{current}})$;
73
74   end foreach
75
76 $\bar{R}_{\text{avail}}\leftarrow(\bar{f}_{1},\bar{f}_{2},\bar{f}_{3},\bar{f}_{4})$;
77
78 $e_{c,k}\leftarrow e_{c,k}\oplus(\bar{R}_{\text{avail}})$;
79
80 end foreach
Algorithm 1 Enhancing an activity instance trace $\sigma_{c}$
3.1.2 Basic (Intra-Case) Features
@@
We extract features from each trace prefix of a trace $\sigma_{c}$. Given a trace $\sigma_{c}=\langle e_{c,1},\dots,e_{c,n}\rangle$ capturing the execution of case $c$, and given an index $1\leq k\leq n$, the trace prefix of case $c$ at index $k$, denoted $\text{$\mathcal{P}$}(c,k)$, is the sequence $\langle e_{c,1},\dots,e_{c,k}\rangle$. For example, in Table 1, the prefix of case 1 at index 2 $\text{$\mathcal{P}$}(1,2)=\langle(\text{Received Query},\text{System},08{:}00,08{:}05),(\text{Assigned to Rep},\text{Supervisor},08{:}10,08{:}15)\rangle$.
@@
For each activity instance $e_{c,k}$ = $(c,a_{c,k},r_{c,k},T^{\text{start}}_{c,k},T^{\text{end}}_{c,k})$ in a prefix, we construct an enhanced activity instance $(c,a_{c,k},r_{c,k},T^{\text{start}}_{c,k},T^{\text{end}}_{c,k},\delta^{\text{start}}_{c,k},\delta^{\text{proc}}_{c,k})$, where $\delta^{\text{start}}_{c,k}=T^{\text{start}}_{c,k}-T^{\text{start}}_{c,k-1}$ denotes the inter-start time, and $\delta^{\text{proc}}_{c,k}=T^{\text{end}}_{c,k}-T^{\text{start}}_{c,k}$ is the processing time. For the first activity, we define $\delta^{\text{start}}_{c,1}=0$. The procedure for extracting vector $v$ is implemented in lines (1-8) of Alg. 1. The algorithm iterates over each trace $\sigma_{c}\in\mathcal{L}$, where $\mathcal{L}$ is the activity instance log grouped by case. For each trace prefix $\text{$\mathcal{P}$}\subseteq\sigma_{c}$, it processes every activity instance, computes the associated intra-case features, and concatenates these features with the tuple representing activity instance $\alpha_{c,i}$ with these features (line 9). [^3] Operator $\oplus$ denotes tuple concatenation.
3.1.3 Resource Contention Features
On top of the above basic features, we add to the feature vectors a set of features to capture resource contention, i.e. the level of busyness of resources. Resource contention is a determinant of waiting times in resource-constrained systems. Accordingly, we expect that these features will enhance the accuracy of inter-start-time predictors.
Prior research [13] highlights the importance of including work in progress (WIP) and process load (e.g., resource utilization) in the context of next-activity and remaining time prediction. Unlike the basic features introduced above, which are extracted from a single trace at a time (intra-case calculation), the WIP is a resource contention feature. Its calculation requires reasoning across multiple cases.
To capture resource contention we propose to capture resource contention by means of three features: WIP, resource utilization, and the arrival rate of new cases. The reason for including WIP and the arrival rate is because of the Little’s Law [9]. According to Little’s Law, the average cycle time ($T_{c}$) of a process is determined by the Work in Progress (WIP) and arrival rate ($\lambda$) via $WIP=\lambda\times T_{c}$. In human-centric processes, cycle time is often dominated by waiting time. We, therefore, hypothesize that features correlated with cycle time can improve waiting time prediction. Queuing theory suggests that waiting times are also influenced by resource utilization, the ratio of resource demand to availability [3]. When demand exceeds capacity, utilization reaches 100%, causing delays, whereas Lower demand reduces utilization. Since our objective is not to predict resources, the task is simplified to predicting the activity instance within a case suffix. Accordingly, we leverage all available and relevant information to effectively predict the case suffix without incorporating resource attributes from the activity instance log.
Alg. 1 explains the extraction of resource contention features, denoted as ${R}_{\text{cont}}$, for a given $\mathcal{P}$. It computes three resource contention features: WIP, resource utilization, and the arrival rate of new cases (lines 10-29). To compute WIP [^4] Throughout this paper, WIP refers to the number of cases currently in execution at a given point in time, i.e., cases that are actively being processed. at time $T^{\text{start}}_{c,k}$, Alg. 1 counts active activity instances in $\mathcal{L}{}$ (lines 11-17). It also estimates the arrival rate as $\lambda=\frac{\text{Total cases}}{\text{Time window}}$ (lines 22-28), using a recent time window (default: 20% of log duration). The set of recent cases is updated dynamically (line 27). Higher WIP and $\lambda$ indicate system congestion, leading to longer expected waiting times. Alg. 1 computes utilization by counting active instances per activity at a given time (lines 16-20), helping predictive models account for resource load and capacity. For each trace prefix $\text{$\mathcal{P}$}\subseteq\sigma_{c}$, Alg. 1 computes resource contention features and augments $e_{c,k}$ accordingly. In addition, Algorithm 1 computes resource availability features using precomputed resource availability calendars, capturing temporal aspects of resource availability; these features are described and discussed in detail in the following section.
3.1.4 Resource Availability Features
In the context of business process simulation, calendars are used to model the availability of resources [16, 18]. A resource availability calendar consists of a set of time intervals during which a resource is on-duty, i.e. available to perform activity instances in a process [21]. Outside these time intervals, the resource is considered to be off-duty, and thus not available to start an activity instance.
Since resources in a business process tend to operate according to daily or weekly cycles (herein called circadian cycles), it is convenient to model an availability calendar by referring to time intervals within a day (e.g. 10:00-11:$0$) and to days within a week (e.g. Mondays). This leads us to the notion ofcircadian time-slot (or time-slot for short). For example, the time-slot “Mondays 10:00-11:00” describes a set of time intervals. Each of these time-intervals consists of a set of time points such that the day of the week is Monday, and the time of the day is between 10:$0$ and 11:$0$.
In line with previous work on resource availability calendars [20], we use quarter-hour (15-minute periods) as the time granularity for availability calendars. In other words, we decompose a day into quarter-hours, e.g. 00:00-00:15, 00:15-00:30 and so on until 23:45-24:$0$ (95 granules per day). This is a sufficiently small granularity to capture typical work shifts, but not too small to potentially overfit. Indeed, if we modelled availability at the granularity of the minute or lower, we could end up with calendars containing time-slots such as Mondays 00:00:00-00:00:20, which would mean that a resource is on-duty for 20 seconds at the start of the day, and then goes off-duty – a situation that is not typical for resources in a business process.
With this convention, we formally define a resource availability calendar as follows.
Definition 3.3 (Resource Availability Calendar).
A circadian time instant $t$ is a tuple $(day,sec)$, where $day\in[0..6]$ denotes the day of the week (Sunday = 0, Saturday = 6), and $sec\in[0,86400]$ represents the number of seconds elapsed since the start of that day.
A time slot($\tau$) is a tuple $(day,start,end)$, where $day\in[0..6]$ specifies the day of the week, and $start,end\in[0..95]$ [^5] Note that there are 96 quarter-hours on a day, herein denoted with numbers 0 to 95. denote the start and end times in units of 15-minute intervals since the beginning of the day. Each interval corresponds to $900$ seconds denoting the number of seconds in a quarter of an hour. The start and end times are such that $start\leq end$.
A calendar is a set of non-overlapping time slots $\{(d_{1},s_{1},e_{1}),(d_{2},s_{2},e_{2}),\dots\}$.
Given a resource $R$, the availability calendar $A(R)$ of $R$ is a calendar consisting of the time-slots when resource $R$ is available to perform work (i.e. to start an activity instance).
The following boolean function allows us to determine if a resource is available (on duty) at a given circadian time point, given the resource’s availability calendar.
[…]
The proposed approach for predicting case suffixes includes resource availability features. These features are extracted from the availability calendars of the resources that are referenced in the event log. Specifically, for each resource, we discover a resource availability calendar using the method reported in [20]. Using the discovered resource availability calendar $A(R)$ for a resource $R$ and given a cutoff point $t_{\text{current}}$, we extract a feature vector, herein referred to as ${R}_{\text{avail}}$, consisting of four features: Time Until Next On-Duty, Time Until Next Off-Duty, Time Since Last On-Duty, and Time Since Last Off-Duty. These features are formally defined below.
Definition 3.4 (Time Until Next On-Duty).
Given a cut-off point $t_{\text{current}}$ and a resource $R$, the Time Until Next On-Duty is the time remaining from $t_{\text{current}}$ until resource $R$ next becomes on duty, taking into consideration the availability calendar of this resource. If the resource is already on duty as of time $t_{\text{current}}$, the Time Until Next On-Duty is 0. Otherwise, it is the time difference between the current time and the next scheduled on-duty time for the resource.
To compute Time Until Next On-Duty, we use two functions DayOfWeek and TimeOfDay, which given an absolute time point, return (respectively) the day of the week and the time since the start of the day, measured in seconds. For example DayOfWeek(‘2026-01-05 00:00:25’) = 6 (since ‘2026-01-10’ is a Saturday) and TimeOfDay(‘2026-01-05 00:00:25’) = 25, since there are 25 elapsed seconds from 00:00:$0$ to 00:00:25.
Given a time-slot $\tau=(day,start,end)$ of a calendar, let DayOfWeek($\tau$) = $\mathit{day}$, i.e. DayOfWeek($\tau$) be the day of the week of time-slot $\tau$. Further, let TimeOfStartDay($\tau$) = $15\times 60\times\textit{(}start)$ be the number of seconds since midnight and the start of time-slot $\tau$. Note that start is the number of quarter-hours since the start of the day and the start of the time-slot. Note also that the factor $15\times 60$ captures the number of seconds in one quarter-hour.
Now, given a time point $t_{\text{current}}$, we define a function $W$ that calculates the time difference (in seconds), between $t_{\text{current}}$ and the start of the week when $t_{\text{current}}$ occurs.
$W(t_{\text{current}})=24\times 60\times 60\times DayOfWeek(t_{\text{current}})+TimeOfDay(t_{\text{current}})$
Note that $24\times 60\times 60=86400$ is the number of seconds in a day.
Similarly, given a time slot $\tau$, we define $W_{s}(\tau)$ as the time elapsed between the start of the week and the start of $\tau$.
$W_{s}(\tau)=24\times 60\times 60\times DayOfWeek(\tau)+TimeOfStartDay(\tau)$
Given a time-slot having start timestamp $\tau$, we can now define the time elapsed from $t_{\text{current}}$ to $\tau_{s}$ as follows:
$TimeElapsed(t_{\text{current}},\tau)=(W_{s}(\tau)-W(t_{\text{current}})+604800)mod604800$
Note that $604800$ is the number of seconds in a week. We add this term to the difference between $t_{\text{current}}$ and $\tau$ and then compute the modulo of the result to account for the fact that circadian time is circular, e.g. the elapsed time from Saturday at 23:59:59 and Sunday at 00:00:01 is not 1604799=6047981-604799=-604798 but rather $(1-604799+604800)\;\mathit{mod}\;604800=2\;\mathit{mod}\;604800=2$.
Given an availability calendar A(R), consisting of a set of timeslots, we define the time until the next time slot as the minimum $\mathit{TimeElapsed}$ from $t_{\text{current}}$ to any time-slot $\tau\in A(R)$. However, we need to take into account that if the resource is on-duty as of time $t_{\text{current}}$, then the time until next on-duty is zero. These observations lead us to the following formula:
[…]
For instance, if a resource is scheduled to start activity instance at 3:$0$ PM (i.e., $\tau^{\text{next}}=15:00:00$) and the current time is 2:30 PM (i.e., $t_{\text{current}}=14:30:00$), the Time Until Next On-Duty will be 30 minutes, or 1800 seconds. On the other hand, if the current time is 4:30 PM and the resource is already on duty, the Time Until Next On-Duty will return 0.
Definition 3.5 (Time Until Next Off-Duty).
Given a cut-off point $t_{\text{current}}$ and a resource $R$, the Time Until Next Off-Duty is the time remaining from $t_{\text{current}}$ until resource $R$ next becomes off duty, taking into consideration the availability calendar of this resource. If the resource is already off duty as of time $t_{\text{current}}$, the Time Until Next Off-Duty is 0. Otherwise, it is the time difference between the current time and the scheduled end of the availability interval in which the resource is currently on duty.
As in Definition 2, we rely on the functions DayOfWeek and TimeOfDay to decompose absolute time points, and we use the function $W$ to map time points to seconds elapsed since the start of the week.
Given a time-slot $\tau=(day,start,end)$ of an availability calendar, let DayOfWeek($\tau$) =day=day, and let
$\text{TimeOfEndDay}(\tau)=15\times 60\times\textit{end},$
which represents the number of seconds since midnight corresponding to the end of the time-slot. We then define
$W_{e}(\tau)=24\times 60\times 60\times\text{DayOfWeek}(\tau)+\text{TimeOfEndDay}(\tau),$
i.e., the number of seconds elapsed between the start of the week and the end of time-slot $\tau$.
The time elapsed from $t_{\text{current}}$ to the end of a time-slot $\tau$ is defined analogously as
$\text{TimeElapsed}(t_{\text{current}},\tau)=\bigl(W_{e}(\tau)-W(t_{\text{current}})+604800\bigr)\bmod 604800,$
where $604800$ is the number of seconds in a week.
Let $A(R)$ denote the availability calendar of resource $R$, and let $A(R,t_{\text{current}})$ indicate whether the resource is on duty at time $t_{\text{current}}$. The Time Until Next Off-Duty is then defined as:
$T_{\text{next\_off\_duty}}(R,t_{\text{current}})=\\ \begin{cases}0,&\text{if }A(R,t_{\text{current}})\text{ is false},\\ \displaystyle\min\{\mathrm{TimeElapsed}(t_{\text{current}},\tau)\\ \mid\tau\in A(R)\},&\text{otherwise.}\end{cases}$ (3)
For instance, if a resource is scheduled to remain on duty until 6:$0$ PM (i.e., the end of the current availability interval is $\tau=18{:}00{:}00$) and the current time is 4:30 PM (i.e., $t_{\text{current}}=16{:}30{:}00$), the Time Until Next Off-Duty is 1.5 hours, or 5400 seconds. On the other hand, if the current time is 7:$0$ PM and the resource is already off duty, the Time Until Next Off-Duty returns 0.
Definition 3.6 (Time Since Last On-Duty).
Given a cut-off point $t_{\text{current}}$ and a resource $R$, the Time Since Last On-Duty measures the time that has elapsed since resource $R$ last became on duty, taking into consideration the availability calendar of the resource. If the resource is currently off duty as of time $t_{\text{current}}$, the Time Since Last On-Duty is 0. Otherwise, it is the time difference between the current time and the start of the availability interval in which the resource is currently on duty.
As in the previous definitions, we rely on the functions DayOfWeek and TimeOfDay to decompose absolute time points, and we use the function $W$ to map time points to seconds elapsed since the start of the week.
Given a time-slot $\tau=(day,start,end)$ of an availability calendar, let DayOfWeek($\tau$) =day=day, and let
$\text{TimeOfStartDay}(\tau)=15\times 60\times\textit{start},$
which represents the number of seconds since midnight corresponding to the start of the time-slot. We then define
$W_{s}(\tau)=24\times 60\times 60\times\text{DayOfWeek}(\tau)+\text{TimeOfStartDay}(\tau_{s}),$
i.e., the number of seconds elapsed between the start of the week and the start of time-slot $\tau$.
The time elapsed from the start of a time-slot $\tau$ to $t_{\text{current}}$ is defined as
$\text{TimeElapsed}(\tau,t_{\text{current}})=\bigl(W(t_{\text{current}})-W_{s}(\tau)+604800\bigr)\bmod 604800,$
where $604800$ is the number of seconds in a week.
Let $A(R)$ denote the availability calendar of resource $R$, and let $A(R,t_{\text{current}})$ indicate whether the resource is on duty at time $t_{\text{current}}$. The Time Since Last On-Duty is then defined as:
$T_{\text{since\_last\_on\_duty}}(R,t_{\text{current}})=\\ \begin{cases}0,&\text{if }A(R,t_{\text{current}})\text{ is false},\\ \displaystyle\min\{\mathrm{TimeElapsed}(\tau,t_{\text{current}})\\ \mid\tau\in A(R)\},&\text{otherwise.}\end{cases}$ (4)
For instance, if a resource started its current availability interval at 1:$0$ PM (i.e., $\tau=13{:}00{:}00$) and the current time is 4:$0$ PM (i.e., $t_{\text{current}}=16{:}00{:}00$), the Time Since Last On-Duty is 3 hours, or 10,800 seconds. If the current time falls outside all availability intervals, the value returned is 0.
Definition 3.7 (Time Since Last Off-Duty).
Given a cut-off point $t_{\text{current}}$ and a resource $R$, the Time Since Last Off-Duty measures the time that has elapsed since resource $R$ last became off duty, taking into consideration the availability calendar of the resource. If the resource is currently on duty as of time $t_{\text{current}}$, the Time Since Last Off-Duty is 0. Otherwise, it is the time difference between the current time and the end of the availability interval immediately preceding the current off-duty period.
As in the previous definitions, we rely on the functions DayOfWeek and TimeOfDay to decompose absolute time points, and we use the function $W$ to map time points to seconds elapsed since the start of the week.
Given a time-slot $\tau=(day,start,end)$ of an availability calendar, let DayOfWeek($\tau$) =day=day, and let
$\text{TimeOfEndDay}(\tau)=15\times 60\times\textit{end},$
which represents the number of seconds since midnight corresponding to the end of the time-slot. We then define
$W_{e}(\tau)=24\times 60\times 60\times\text{DayOfWeek}(\tau)+\text{TimeOfEndDay}(\tau),$
i.e., the number of seconds elapsed between the start of the week and the end of time-slot $\tau$.
The time elapsed from the end of a time-slot $\tau$ to $t_{\text{current}}$ is defined as
$\text{TimeElapsed}(\tau,t_{\text{current}})=\bigl(W(t_{\text{current}})-W_{e}(\tau)+604800\bigr)\bmod 604800,$
where $604800$ is the number of seconds in a week.
Let $A(R)$ denote the availability calendar of resource $R$, and let $A(R,t_{\text{current}})$ indicate whether the resource is on duty at time $t_{\text{current}}$. The Time Since Last Off-Duty is then defined as:
$T_{\text{since\_last\_off\_duty}}(R,t_{\text{current}})=\\ \begin{cases}0,&\text{if }A(R,t_{\text{current}})\text{ is true},\\ \displaystyle\min\{\mathrm{TimeElapsed}(\tau,t_{\text{current}})\\ \mid\tau\in A(R)\},&\text{otherwise.}\end{cases}$ (5)
For instance, if a resource ended its most recent availability interval at 5:$0$ PM (i.e., $\tau=17{:}00{:}00$) and the current time is 7:$0$ PM (i.e., $t_{\text{current}}=19{:}00{:}00$), the Time Since Last Off-Duty is 2 hours, or 7,200 seconds. If the resource is currently on duty, the value returned is 0.
Algorithm 1 computes resource availability features by querying precomputed resource availability calendars $C_{r}$ obtained following [20]. For each trace prefix $\text{$\mathcal{P}$}\subseteq\sigma_{c}$, availability is evaluated at the current decision time $T^{\text{current}}$, defined as the start time of the active activity instance $e_{c,k}$.The algorithm first identifies the set of resources $\mathcal{R}^{a_{k}}=\mathcal{R}(a_{c,k})$ that can execute the activity $a_{c,k}$ (line 30). For each eligible resource $r\in\mathcal{R}^{a_{k}}$, four temporal availability features are extracted from the corresponding calendar $C_{r}$ at time $T^{\text{current}}$: time to next on-duty, time since last on-duty, time to next off-duty, and time since last off-duty (lines 31--35). These features are aggregated across all eligible resources by computing the minimum value per feature dimension (lines 36--38), yielding an availability vector $\bar{R}_{\text{avail}}$. Finally, this vector is appended to the active activity instance $e_{c,k}$, enriching it with resource availability information used by case suffix prediction models.
3.1.5 Feature Vector Construction